# REFRAME: Reflective Surface Real-Time Rendering for Mobile Devices

### v1

@@ Abstract
This work tackles the challenging task of achieving real-time novel view synthesis on various scenes, including highly reflective objects and unbounded outdoor scenes. Existing real-time rendering methods, especially those based on meshes, often have subpar performance in modeling surfaces with rich view-dependent appearances. Our key idea lies in leveraging meshes for rendering acceleration while incorporating a novel approach to parameterize view-dependent information. We decompose the color into diffuse and specular, and model the specular color in the reflected direction based on a neural environment map. Our experiments demonstrate that our method achieves comparable reconstruction quality for highly reflective surfaces compared to state-of-the-art offline methods, while also efficiently enabling real-time rendering on edge devices such as smartphones. Our project page is at https://xdimlab.github.io/REFRAME/.
@@ 1 Introduction
There are many existing methods focusing on accelerating the rendering of NeRF. One line of works sticks with the volume rendering pipeline where the color of a pixel is accumulated along the ray. Acceleration is achieved by improving the sampling strategy [47, 25, 31, 27], tabulating the intermediate output [17, 12, 56, 48, 15, 53, 40, 34], or leveraging super-resolution neural rendering [28]. However, deploying these methods on edge devices such as smartphones is often greatly hindered due to the demanding computational power they typically require. Besides, tabulation-based methods require large memory consumption and greatly increase communication costs when transmitting data between the cloud server and the client. Another line of works [45, 39, 9] distills radiance fields into a mesh for real-time rendering, combined with a small MLP to model view-dependent effects. Mesh-based methods can leverage traditional graphics pipelines for acceleration, enabling them to achieve real-time rendering even on edge devices. Nonetheless, real-time rendering methods [45, 9, 15, 56] often struggle to model objects with highly reflective surfaces. Besides, these mesh-based real-time rendering methods [55, 9, 45] typically require a large number of vertices and faces to achieve high fidelity.
@@ 2 Related Work
Reflectance Decomposition: A number of works [50, 58, 57, 5, 4, 59, 43, 61, 62, 19] investigate the task of inferring geometry and material properties based on neural fields representation, typically by formulating the image generation using the physically-based rendering. Despite achieving lighting and material control, these methods are typically inferior to state-of-the-art NeRF-based methods that directly model the radiance in terms of the rendering quality. This is due to the fact that these methods rely on simplified rendering equations of the real world. Another line of works achieves better rendering quality by avoiding full decomposition, yet allowing for modeling glossy objects by modeling the reflected radiance [46, 14, 54] or replacing the explicit rendering equation with a learned neural renderer [30, 6]. However, none of these methods are capable of real-time rendering due to the underlying volume rendering formulation. While NvDiffRec [35] enables real-time rendering based on the mesh representation, its quality is also limited by the full decomposition. In this work, we follow the partial decomposition pipeline [46] and propose to learn a neural environment map based on a NeRF-distilled mesh representation, with a focus on achieving real-time rendering.
@@ 3.1 Geometry Learner
Before training, we follow NeRF2Mesh [45] and use grid representation [34] to extract our initial coarse mesh. Then, we propose to leverage a geometry learner to refine the geometry of the initial coarse mesh, including refining the vertex positions and normals.
Vertex Offset: As shown in Fig. 2, the initial mesh extracted from NeRF-based methods may be of low quality. Existing methods [45, 52] directly update the vertex positions through gradient backpropagation. However, our experimental results have shown that this update strategy lacks robustness. The step size of the update is determined by the learning rate and the gradients backpropagated, where a large learning rate can lead to catastrophic consequences while a small learning rate results in slow convergence. Hence, we introduce a network to directly learn vertex updates:
@@
where $\mathbf{v}^{\prime}$ is the updated vertex, $g_{\theta_{\mathbf{v}}}$ is a multi-resolution hash encoding in combined with a small MLP [34]. Utilizing a multi-resolution hash grid allows us to leverage local information as well as global information across different positions, while the direct gradient-based update method can only utilize local information.
@@ 3.2 Color Formulation
where $\mathbf{c}_{d}+\mathbf{c}_{s}$ is clamped to $[0,1]$. We now elaborate on the diffuse and specular color formulation.
@@ 3.3 Loss Functions
We also introduce to use SSIM [16] loss for training. This enhances the perceptual quality of our rendered results.
On object-centric datasets, we additionally utilize mask loss to ensure the rasterization mask closely aligns with the ground truth mask. Note that the mask loss is optional and is not required on unbounded datasets.
@@ 3.4 Real-Time Rendering
Texture Images: To further reduce the inference burden and better deploy our method to mobile devices or the web platform on desktop, we can also bake the $\mathbf{c}_{d}$ and $\mathbf{f}_{s}$ obtained from querying the $f_{\phi}$ into two texture images. Specifically, we map the vertices of the mesh to UV coordinates and then bake the features into two texture images, respectively. In order to obtain the normal $\mathbf{n}_{\mathbf{p}}$ from the learned normal vector $\mathbf{n}^{\prime}_{\mathbf{v}}$ during real-time rendering, we also bake the normal $\mathbf{n}_{\mathbf{p}}$ into a texture image.
@@ 4 Experiment
Dataset. We validate the effectiveness and robustness of our method using three different datasets with varied scenes. 1) NeRF [33] Synthetic Dataset: This dataset consists of eight synthetic scenes. 2) Shiny Blender Dataset: This dataset is introduced by Ref-NeRF [46] and includes six glossy objects, serving as the primary dataset in the experimental section for assessing the reflective surface modeling capabilities of the methods. 3) Real Captured Dataset: We select three outdoor scenes with rich reflective appearances same as Ref-NeRF [46] to validate the effectiveness of our method in real-world outdoor capture scenarios. Due to the poor quality of the initial mesh extracted by Nerf2Mesh [45] when modeling glossy objects, we utilize the meshes extracted by Neuralangelo [29] and Ref-NeuS [14] as our initialization in certain scenes, see supplementary for more details.
@@ 4.2 Rendering Efficiency
Table 2: Computational Efficiency Comparison. We test the FPS of our method and the baseline methods [45, 9] on the NeRF Synthetic dataset across multiple devices. On MacBook Pro, iPhone 12, and iPad Air3, both our method and the baseline methods [45, 9] achieve the maximum screen refresh rate.
@@
Memory efficiency. Our cache mainly consists of texture images, the environment feature map, and the mesh. Due to our high-quality estimation of vertex normals, our mesh can achieve high-quality rendering while having a lower number of vertices and faces, leading to a low cache size in most object-level scenes, as shown in Tab. 3.
@@
Table 3: Memory Efficiency. While exhibiting higher rendering quality, our method maintains comparable or even lower memory consumption compared to the other methods. Cache reported in MB.
@@ 4.3 Ablation Study
In order to verify the necessity of estimating an offset for the vertices normal, we conduct experiments with the w/o normal offset setting. In this setting, we do not estimate a normal offset for each vertex, but we still compute an updated normal based on the updated vertex positions. As shown in Tab. 5, introducing per-vertex normal offset provides a gain in reconstruction quality across multiple datasets. This gain is particularly significant when the number of faces is low. Notably, in the case of the toaster dataset, a mesh with 10,000 faces with normal offset even achieves rendering quality surpassing that of a mesh with 75,000 faces without normal offset.
@@
Refer to caption (a) w/o $f_{e}$ Refer to caption (b) Ours Refer to caption (c) Ground Truth
w/o $f_{e}$ Ours Toaster 19.31 24.95 Materials 27.37 29.58 Table 6: Ablation of Environment Learner. Reported in PSNR.
@@ A.3 Training Implementation
Training Details: Our method utilizes the Adam [23] optimizer and employs the cosine annealing learning rate adjustment strategy [32] with a period of 2 and a period step of 2. We train our model for 250 epochs on each scene. With a batch size of 1, we render only one image per iteration and calculate the loss to optimize our network. During training, we set $\lambda_{\mathbf{c}}$ to 1, $\lambda_{\mathbf{c}_{d}}$ to 0.001, $\lambda_{SSIM}$ to 3, $\lambda_{mask}$ to 100, $\lambda_{\Delta\mathbf{n}_{\mathbf{v}}}$ to 0.1, $\lambda_{max}$ to 0.00001.
@@ C.1 Reconstruction Quality
Table 8: PSNRnormal-$\uparrow$ on NeRF Synthetic Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 9: SSIMnormal-$\uparrow$ on NeRF Synthetic Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 10: LPIPSnormal-$\downarrow$ on NeRF Synthetic Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 11: PSNRnormal-$\uparrow$ on Shiny Blender Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 12: SSIMnormal-$\uparrow$ on Shiny Blender Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 13: LPIPSnormal-$\downarrow$ on Shiny Blender Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 14: PSNRnormal-$\uparrow$ on Real Captured Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 15: SSIMnormal-$\uparrow$ on Real Captured Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.
@@
Table 16: LPIPSnormal-$\downarrow$ on Real Captured Dataset. Red represents the optimal, orange represents the second best, and yellow represents the third.

### v2

@@ Abstract
This work tackles the challenging task of achieving real-time novel view synthesis for reflective surfaces across various scenes. Existing real-time rendering methods, especially those based on meshes, often have subpar performance in modeling surfaces with rich view-dependent appearances. Our key idea lies in leveraging meshes for rendering acceleration while incorporating a novel approach to parameterize view-dependent information. We decompose the color into diffuse and specular, and model the specular color in the reflected direction based on a neural environment map. Our experiments demonstrate that our method achieves comparable reconstruction quality for highly reflective surfaces compared to state-of-the-art offline methods, while also efficiently enabling real-time rendering on edge devices such as smartphones. Our project page is at https://xdimlab.github.io/REFRAME/.
@@ 1 Introduction
There are many existing methods focusing on accelerating the rendering of NeRF. One line of works sticks with the volume rendering pipeline where the color of a pixel is accumulated along the ray. Acceleration is achieved by improving the sampling strategy [49, 26, 32, 28], tabulating the intermediate output [18, 12, 59, 50, 16, 56, 41, 35], or leveraging super-resolution neural rendering [29]. However, deploying these methods on edge devices such as smartphones is often greatly hindered due to the demanding computational power they typically require. Besides, tabulation-based methods require large memory consumption and greatly increase communication costs when transmitting data between the cloud server and the client. Another line of works [46, 40, 9] distills radiance fields into a mesh for real-time rendering, combined with a small MLP to model view-dependent effects. Mesh-based methods can leverage traditional graphics pipelines for acceleration, enabling them to achieve real-time rendering even on edge devices. Nonetheless, real-time rendering methods [46, 9, 16, 59, 15] often struggle to model objects with highly reflective surfaces. Besides, these mesh-based real-time rendering methods [58, 9, 46] typically require a large number of vertices and faces to achieve high fidelity.
@@ 2 Related Work
Reflectance Decomposition: A number of works [53, 62, 61, 5, 4, 63, 44, 65, 66, 20, 60] investigate the task of inferring geometry and material properties based on neural fields representation, typically by formulating the image generation using the physically-based rendering. Despite achieving lighting and material control, these methods are typically inferior to state-of-the-art NeRF-based methods that directly model the radiance in terms of the rendering quality. This is due to the fact that these methods rely on simplified rendering equations of the real world. Another line of works achieves better rendering quality by avoiding full decomposition, yet allowing for modeling glossy objects by modeling the reflected radiance [47, 14, 57] or replacing the explicit rendering equation with a learned neural renderer [31, 6]. However, none of these methods are capable of real-time rendering due to the underlying volume rendering formulation. While NvDiffRec [36] enables real-time rendering based on the mesh representation, its quality is also limited by the full decomposition. In this work, we follow the partial decomposition pipeline [47] and propose to learn a neural environment map based on a NeRF-distilled mesh representation, with a focus on achieving real-time rendering.
@@ 3.1 Geometry Learner
Before training, we follow NeRF2Mesh [46] to extract the initial coarse mesh through grid-based representation [35]. This allows for better modeling of depth discontinuities in geometry [51] and prevents the optimization of mesh geometry from getting stuck in local optima. Then, we propose to leverage a geometry learner to refine the geometry of the initial coarse mesh, including refining the vertex positions and normals.
Vertex Offset: As shown in Fig. 2, the initial mesh extracted from NeRF-based methods may be of low quality. Existing methods [46, 55] directly update the vertex positions through gradient backpropagation, which means the vertex update step size is determined by the learning rate and the gradient backpropagated. Thus, costly and difficult manual learning rate tuning is needed to adapt update step sizes for different objects, with a lack of robustness as shown in our experiments. To address this, our method uses a hybrid representation to provide a learned, adaptive step size for different vertices and different objects:
@@
where $\mathbf{v}^{\prime}$ is the updated vertex, $g_{\theta_{\mathbf{v}}}$ is a multi-resolution hash encoding in combined with a small MLP [35]. Though [48] also uses MLP for vertex offset learning, our coarse-to-fine hybrid representation allows us to leverage local information as well as global information, leading to better modeling of the geometry.
@@ 3.2 Color Formulation
where $\mathbf{c}_{d}+\mathbf{c}_{s}$ is clamped to $[0,1]$, both individual terms $\mathbf{c}_{d}$ and $\mathbf{c}_{s}$ from sigmoid are in [0,1]. We now elaborate on the diffuse and specular color formulation.
@@ 3.3 Loss Functions
We also introduce SSIM [17] loss for better perceptual quality. On object-centric datasets, we additionally utilize mask loss to ensure the rasterization mask closely aligns with the ground truth mask. Note that the mask loss is optional and is not required on unbounded datasets.
@@ 3.4 Real-Time Rendering
Texture Images: To further reduce the inference burden and better deploy our method across different platforms, we can also bake the $\mathbf{c}_{d}$ and $\mathbf{f}_{s}$ obtained from querying the $f_{\phi}$ into two texture images. Specifically, we map the vertices of the mesh to UV coordinates and then bake the features into two texture images, respectively. In order to obtain the normal $\mathbf{n}_{\mathbf{p}}$ from the learned normal vector $\mathbf{n}^{\prime}_{\mathbf{v}}$ during real-time rendering, we also bake the normal $\mathbf{n}_{\mathbf{p}}$ into a texture image.
@@ 4 Experiment
Dataset. We validate the effectiveness and robustness of our method using three different datasets with varied scenes. 1) NeRF [34] Synthetic Dataset: This dataset consists of eight synthetic scenes. 2) Shiny Blender Dataset: This dataset is introduced by Ref-NeRF [47] and includes six glossy objects, serving as the primary dataset for assessing the reflective surface modeling capabilities of the methods. 3) Real Captured Dataset: We select three outdoor scenes with rich reflective appearances same as Ref-NeRF [47] to validate the effectiveness of our method in real-world outdoor capture scenarios. Due to the poor quality of the initial mesh extracted by Nerf2Mesh [46] when modeling glossy objects, we utilize the meshes extracted by Neuralangelo [30] and Ref-NeuS [14] as our initialization in certain scenes, see supplementary for more details.
@@ 4.2 Rendering Efficiency
Table 2: Computational Efficiency Comparison. We test the FPS of our method and the baseline methods [46, 9] on the NeRF Synthetic dataset across multiple devices.
@@
Memory efficiency. Our cache mainly consists of texture images, the environment feature map, and the mesh. Due to our high-quality estimation of vertex normals, our mesh can achieve high-quality rendering with a low number of vertices and faces, leading to a small cache size in most scenes, as shown in Tab. 3.
@@
Table 3: Memory Efficiency. While exhibiting higher quality, our method maintains a relatively low memory consumption compared to baselines. Cache reported in MB.
@@ 4.3 Ablation Study
In order to verify the necessity of estimating an offset for the vertices normal, we conduct experiments with the w/o normal offset setting. In this setting, we do not estimate a normal offset for each vertex, but we still compute an updated normal based on the updated vertex positions. As shown in Tab. 5, introducing per-vertex normal offset provides a gain in reconstruction quality. This gain is particularly significant when the number of faces is low. Notably, in the case of the toaster dataset, a mesh with 10k faces with normal offset even achieves rendering quality surpassing that of a mesh with 75k faces without normal offset.
@@
w/o $f_{e}$ Ours
Toaster 19.31 24.95
Materials 27.37 29.58
Table 6: Ablation of Environment Learner. Reported in PSNR.
@@ 5 Conclusion
Acknowledgements
This work is supported by NSFC under grant 62202418 and U21B2004. We thank Xiaoran Cao, Chenxi Tu, and Sicheng Li for their valuable discussions.
@@ A.3 Training Implementation
Training Details: Our method utilizes the Adam [24] optimizer and employs the cosine annealing learning rate adjustment strategy [33] with a period of 2 and a period step of 2. We train our model for 250 epochs on each scene. With a batch size of 1, we render only one image per iteration and calculate the loss to optimize our network. During training, we set $\lambda_{\mathbf{c}}$ to 1, $\lambda_{\mathbf{c}_{d}}$ to 0.001, $\lambda_{SSIM}$ to 3, $\lambda_{mask}$ to 100, $\lambda_{\Delta\mathbf{n}_{\mathbf{v}}}$ to 0.1, $\lambda_{max}$ to 0.00001. Given the coarse initial mesh, we scale normals and vertex offsets with $w=\min(e/e_{warm},1)$, where $e$ represents the current epoch and $e_{warm}$ is the warm-up epoch number to stabilize training process.
@@ C.1 Reconstruction Quality
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]